OpenStudy (anonymous):

Rationalize the numerator of the expression. Then simplify your answer. (√7 – 3)/5. my answer is incorrect. i have -2/5(√7+3)

OpenStudy (anonymous):

√7 – 3) * √7 + 3) = 7 - 9 = -2 also multiply 5 by (√7 + 3): (√7 – 3)/5 = -2 / 5(√7 + 3) - looks like you are right

OpenStudy (anonymous):

it said i was wrong though. should i distribute the 5? because it says to simplify.

OpenStudy (radar):

\[\sqrt{7-3}\over5 \]???

OpenStudy (radar):

2/5

OpenStudy (anonymous):

yea - could be - sqrt refers to (7-3)

OpenStudy (anonymous):

its just √7-3, the √is only above the 7.

OpenStudy (radar):

Never mind

OpenStudy (anonymous):

the question as you typed it is a bit odd - can you check the original?

OpenStudy (anonymous):

(√7-3)/(5)

OpenStudy (anonymous):